# Why In-Context Learning Transformers are Tabular Data Classifiers

### v1

@@ Abstract
The recently introduced TabPFN pretrains an In-Context Learning (ICL) transformer on synthetic data to perform tabular data classification. As synthetic data does not share features or labels with real-world data, the underlying mechanism that contributes to the success of this method remains unclear. This study provides an explanation by demonstrating that ICL-transformers acquire the ability to create complex decision boundaries during pretraining. To validate our claim, we develop a novel forest dataset generator which creates datasets that are unrealistic, but have complex decision boundaries. Our experiments confirm the effectiveness of ICL-transformers pretrained on this data. Furthermore, we create TabForestPFN, the ICL-transformer pretrained on both the original TabPFN synthetic dataset generator and our forest dataset generator. By fine-tuning this model, we reach the current state-of-the-art on tabular data classification. Code is available at https://github.com/FelixdenBreejen/TabForestPFN.
@@ 1 Introduction
Tabular data classification is widespread across all industries, leading to an increased interest in the research field of deep learning for tabular data [28, 52, 25, 35]. This type of classification involves classifying a target variable based on a set of attributes, which is commonly stored in tabular format. Examples of tabular classification include predicting the existence of chronic kidney disease based on blood test results [34], estimating the click-through rate of advertisements [38], and predicting the stability of pillars in hard rock mines [29]. Despite the significance of tabular data, major breakthroughs in AI as shown on vision [16, 19, 9] and language [48, 3] data have yet to reach the tabular domain. In fact, neural networks are currently outperformed by tree-based machine learning algorithms such as XGBoost [6] and CatBoost [37] in tabular classification tasks [12, 15, 32].
@@
In an attempt to bridge this performance gap, a recent method called tabular prior-data fitted networks (TabPFN) [20] introduces an in-context learning (ICL) [8] scheme, demonstrating promising results [15]. This tabular ICL-transformer can predict test observations with only one forward pass using training observations included in the context. Hollmann et al. [20] generate their pretraining data synthetically, with a focus on creating realistic datasets that act as a “prior”. They make their datasets realistic by carefully crafting correlations between features, introducing variety in the feature importance, and by leveraging structural causal models, which can simulate causal relationships.
Even with realistic datasets, it is questionable why pretraining ICL-transformers would improve classification performance on real-world tabular benchmarks. While pretraining in vision and language has clear purposes by transferring texture and grammar knowledge, explanations from a transfer learning point of view in the field of tabular data do not make sense, as datasets do not share features or labels.
@@
In this work, we claim that the effectiveness of ICL-transformers stems from their ability to form complex decision boundaries. See Figure 1 for an example of such a decision boundary. Intuitively, the complexity is given by how far the decision boundary differs from a simple linear line. Neural networks that train from scratch on tabular data often have overly simple decision boundaries, a phenomenon known as simplicity bias [41], while tree-based methods do not suffer from this [15]. We show that ICL-transformers, in contrast, learn how to create these complex decision boundaries during pretraining. To support our claim, we present three arguments:
  • ICL-transformers trained on a highly complex but unrealistic dataset generator achieve performance similar to TabPFN. We create this novel forest dataset generator based on decision trees, and we call the ICL-transformer pretrained on this data TabForest. Although the generator outputs orthogonal decision boundaries and independent and identically distributed features, the performance of TabForest on individual benchmark datasets is strikingly similar to TabPFN.
  • Pretraining dataset complexity is highly correlated with performance. To show this, we vary the parameters of our forest dataset generator that influence the complexity of the decision boundaries. We observe that higher complexity increases the performance.
  • ICL-transformers can create decision boundaries with high complexity. This property emerges when fine-tuning on a specific dataset, as illustrated in Figure 1. Furthermore, we observe that complexity increases with the size of the model.
@@
Figure 1: Comparison of decision boundaries for the Electricity dataset (OpenML ID 44156). Axis represent features, colors are predicted class probabilities, and dots are test observations. Fine-tuned variants show a higher complexity score (see section 5.4) than non-fine-tuned variants (Zero-shot).
In addition to explaining why ICL-transformers work, we push their performance to new heights. TabForestPFN, the ICL-transformer trained on both the TabPFN and the forest dataset, reaches state of the art on two benchmarks [15, 32]. This ensemble boosts performance by seemingly selecting the best method between TabPFN and TabForest, leveraging the small unexplained differences in performance across various real-world datasets.
We achieve a second major push in performance by switching from zero-shot to fine-tuning. The additional fine-tuning phase improves performance on average, even when the number of observations is less than a thousand. Moreover, both fine-tuning and zero-shot performance highly depends on the context size, where larger contexts yield better results, even far beyond the context size pretrained on. These findings suggest that fine-tuning, using the largest possible context size, should be the default choice when working with ICL-transformers.
@@
In conclusion, ICL-transformers are highly effective tabular data classifiers, capable of creating complex decision boundaries. This new insight advances our understanding of tabular ICL-transformers and opens up new avenues for further research to enhance their performance. With further developments, we anticipate a significant shift in the field of tabular data, moving from tree-based methods towards ICL-transformers.
@@ 2 Related Works
There are three main branches of tools for tabular data classification: classical statistical methods like linear regression, K-nearest neighbors, and support vector machines [17]; tree-based algorithms like XGBoost [6], CatBoost [37], and LightGBM [24]; and neural-network based methods like this paper. There are several papers benchmarking the different methods [12, 43, 15, 32, 50]. Overall, tree-based methods stand at the top, with neural networks ranging from inferior to at best competitive.
Nonetheless, there have been numerous approaches that tackle tabular data classification with neural networks. First, we have the class of neural networks trained from scratch: training starts from random initialized weights and is only trained on the data at hand. Research has focused on architectures [23, 44, 1, 14, 21, 4], embeddings [39, 13, 5], and regularization [42, 22].
In general, methods training from scratch can struggle because tabular datasets can be small. So, researchers have sought ways to use large volumes of tabular data or to change the training objective. Some employ self-supervised learning [26, 49, 54, 2, 46, 45], or closely related transfer learning techniques [33, 27, 53]. Others leverage pretrained LLMs [18, 51] to come to their prediction.
One of those related transfer learning methodologies is in-context learning, a new field sparked by TabPFN [20]. Currently, there is ongoing research on how to scale TabPFN to encompass more observations and features [30, 10, 11]. Amid these developments, we present our work.
@@ 3 Preliminaries
In tabular classification, we are interested in predicting targets $y\in{\mathbb{N}}$ given features ${\bm{x}}\in{\mathbb{R}}^{d}$, where $d$ is the number of features. We predict $y$ using an in-context learning (ICL) transformer pretrained on a synthetic dataset. The in-context learning allows the transformer to predict targets based on other observations included in the forward pass. In our work, zero-shot refers to one forward pass through the ICL-transformer without any fine-tuning, while fine-tune refers to one forward pass through the ICL-transformer after fine-tuning. Our work builds on TabPFN [20], so below we explain their dataset generator and their transformer architecture.
@@ 3.1 TabPFN Dataset Generator
The TabPFN authors create their own synthetic dataset using Bayesian Neural Networks (BNN) and Structural Causal Models (SCM). To construct a dataset, they first create a BNN or a SCM with random characteristics and with randomly initialized weights. Then they randomly draw an input ${\bm{X}}$ and pass it through the model to generate output ${\bm{y}}$. Their final dataset is given by $({\bm{X}},{\bm{y}})$. See their paper [20] for more details.
In their approach, they emphasize their ability to create realistic datasets, and even call their generator a "prior". They chose SCMs specifically because it can capture real-world causal mechanisms. One other aspect they focus on is simplicity, biasing the generator towards less complex input-output relationships. Additionally, they ensure the inputs ${\bm{X}}$ have natural correlation by correlating the features blockwise, and they vary their feature importance by tuning the magnitude of weights belonging to different features. These methods suggest the authors believe creating realistic datasets is important for achieving good performance.
@@ 3.2 Architecture
In our work, we use the architecture from TabPFN, and make no changes to isolate the effect of the dataset generator. This ICL-transformer has as input the features ${\bm{X}}_{support}\in{\mathbb{R}}^{|S|\times d_{f}}$ and targets ${\bm{y}}_{support}\in{\mathbb{N}}^{|S|}$ from support set $S$ and features ${\bm{X}}_{query}\in{\mathbb{R}}^{|Q|\times d_{f}}$ from query set $Q$. The output is a prediction for ${\bm{y}}_{query}\in{\mathbb{R}}^{|Q|}$. The query set $Q$ represents the observations we want to predict, while the support set $S$ are the observations we base our prediction on. This architecture accepts a fixed number of features $d_{f}$, see also the preprocessing discussed in Appendix A.3.
In this transformer, a token with dimension $d_{token}$ represents all features of a single observation. The creation of support tokens $H_{support}\in{\mathbb{R}}^{|S|\times d_{token}}$ and query tokens $H_{query}\in{\mathbb{R}}^{|Q|\times d_{token}}$ are given by equations 1 and 2.
@@ 4.1 Forest Dataset Generation
Our goal is to create a simple dataset generator that generates datasets with complex patterns to train on, in contrast to the TabPFN [20] generator that aims to create realistic datasets. We base our generator on decision trees, because of their ability to create highly complex decision boundaries with minimal computation cost. The idea is to "overfit" the decision tree to randomly generated features and targets. This fitted decision tree is then used as a data-generating process. See Algorithm 1 for the method and Figure 3 for examples of generated data.
@@
Our forest dataset generator allows datasets to vary in number of classes, observations, numerical features, and categorical features. There are two parameters that contribute to the decision boundary complexity. The base size is the number of observations used to fit the decision tree; more observations means more places for the decision tree to split on. The tree depth determines how deep the decision tree will go before exiting the fitting algorithm, with higher depth leading to greater complexity.
For every new synthetically generated dataset, we uniformly draw the hyperparameters from the bounds shown in Table 1. Because we have two hyperparameters that influence the complexity, we decided to keep the base size fixed. In the final step of the algorithm, the targets ${\bm{y}}_{2}$ are discretized by uniformly drawing bucket boundaries between 0.0 and 1.0, and assigning a class to each bucket, creating varying degrees of class imbalance.
Table 1: Hyperparameters for the Forest Dataset Generator
Hyperparameter min max
base size 1024 1024
dataset size 128 1024
tree depth 1 25
number of features 3 100
number of classes 2 10
ratio of categorical features 0.0 1.0
Figure 3: Generated forest data. Every box is a generated dataset with its own classes (color) and features (axes). The data clouds look unrealistic: decision boundaries are always orthogonal, and there is no feature correlation. Generated with base size 1024, dataset size 1024, maximum tree depth between 1 and 25, two features, and between 2 and 10 number of classes.
@@ 4.2 Fine-tuning Procedure
In our work, we introduce fine-tuning to the tabular ICL-transformer. When fine-tuning, we like to draw support and query sets from our training data such that the performance generalizes to the test set. This requires careful consideration of dataset splitting. The benchmark datasets already provide us with a training-validation-test split, of which the validation set is used for hyperparameter search. We require an additional validation set for early stopping, so we split the training with ratio 80/20 in training-small and validation-early-stopping.
Every gradient decent step, we randomly draw a 80/20 support and query split from training-small. For validation, we draw the support set from training-small, and draw the query set from validation-early-stopping. Every sample of the validation-early-stopping is seen exactly once, while support samples are randomly drawn with replacement. We use early stopping based on the validation loss, which is calculated after every step.
The early stopping technique can decide that training zero epochs is optimal, which allows us to fall back on the zero-shot performance in case fine-tuning harms the performance. This is especially important when using very small datasets. At the same time, fine-tuning can make use of the full training dataset, while the number of samples zero-shot can use is limited by the maximum support size that fits on the GPU.
@@ 5 Experiments
In our results we consider four pre-trained architectures, each with a zero-shot and a fine-tuned version:
  • TabPFN (original) is the original implementation by the TabPFN [20] authors, fine-tuned by us. The weights are downloaded from their Github.
  • TabPFN (retrained) is trained by us on the TabPFN-dataset.
@@
Comparing the difference in behavior and performance allows us to learn the effect of the different synthetic datasets. Training and hyperparameter settings are given in appendix A.4, benchmarks used and results obtained are given below.

### v2

# Fine-tuned In-Context Learning Transformers are Excellent Tabular Data Classifiers

@@ Abstract
The recently introduced TabPFN pretrains an In-Context Learning (ICL) transformer on synthetic data to perform tabular data classification. In this work, we extend TabPFN to the fine-tuning setting, resulting in a significant performance boost. We also discover that fine-tuning enables ICL-transformers to create complex decision boundaries, a property regular neural networks do not have. Based on this observation, we propose to pretrain ICL-transformers on a new forest dataset generator which creates datasets that are unrealistic, but have complex decision boundaries. TabForest, the ICL-transformer pretrained on this dataset generator, shows better fine-tuning performance when pretrained on more complex datasets. Additionally, TabForest outperforms TabPFN on some real-world datasets when fine-tuning, despite having lower zero-shot performance due to the unrealistic nature of the pretraining datasets. By combining both dataset generators, we create TabForestPFN, an ICL-transformer that achieves excellent fine-tuning performance and good zero-shot performance.
@@ 1 Introduction
Tabular data classification is widespread across all industries, leading to an increased interest in the research field of deep learning for tabular data (Liakos et al., 2018; Zhang et al., 2020; Keith et al., 2021; Pang et al., 2022). This type of classification involves classifying a target variable based on a set of attributes, which are commonly stored in tabular format. Examples of tabular classification include predicting the existence of chronic kidney disease based on blood test results (Ogunleye & Wang, 2020), estimating the click-through rate of advertisements (Richardson et al., 2007), and predicting the stability of pillars in hard rock mines (Liang et al., 2020). Despite the significance of tabular data, major breakthroughs in AI, as demonstrated in vision and language domains, have yet to reach the tabular domain. In fact, neural networks are currently outperformed by tree-based machine learning algorithms such as XGBoost (Chen & Guestrin, 2016) and CatBoost (Prokhorenkova et al., 2018) in tabular classification tasks (Gorishniy et al., 2021; Grinsztajn et al., 2022; McElfresh et al., 2023).
@@
In an attempt to bridge this performance gap, a recent method called tabular prior-data fitted networks (TabPFN) (Hollmann et al., 2023) introduces an in-context learning (ICL) (Dong et al., 2023) scheme, demonstrating promising results (Grinsztajn et al., 2022). This tabular ICL-transformer can predict test observations zero-shot: with only one forward pass using training observations included in the context. Hollmann et al. generate their pretraining data synthetically, focusing on creating realistic datasets that act as a “prior”. They make their datasets realistic by carefully crafting correlations between features, introducing variety in feature importance, and leveraging structural causal models to simulate causal relationships.
In this work, we extend TabPFN to the fine-tuning setting. We show that fine-tuning this ICL-transformer on downstream tasks boosts the performance, particularly on datasets with more than a thousand samples. We also find that increasing the context size provides a performance increase for both zero-shot and fine-tuning. Overall, we find that fine-tuning provides such a large performance boost that we recommend always using it over zero-shot when the number of observations exceeds a thousand, and using zero-shot only when inference speed is an issue.
@@
During this process, we also discovered an interesting property of ICL-transformers. Fine-tuned ICL-transformers can create complex decision boundaries, see Figure 1 for an example. Intuitively, the complexity is determined by how far the decision boundary differs from a simple linear line. Neural networks trained from scratch on tabular data often have overly simple decision boundaries, a phenomenon known as simplicity bias (Shah et al., 2020), while tree-based methods do not suffer from this (Grinsztajn et al., 2022). We find that fine-tuned ICL-transformers, in contrast, are able to create these complex decision boundaries similar to tree-based methods.
Given this observation, we wonder whether pretraining on more complex data would improve the fine-tuning performance. To this end, we introduce the novel forest dataset generator, which creates highly complex synthetic datasets using a simple decision tree. By varying the parameters of the decision tree, we can control the complexity of the generated datasets. We observe that TabForest, the ICL-transformer pretrained on this forest dataset generator, achieves better fine-tuning performance as the complexity of the generated datasets increases.
Furthermore, TabForest shows fine-tuning performance that surpasses TabPFN on specific real-world datasets, even though the zero-shot performance of TabForest is significantly lacking compared to TabPFN. This suggests that for the fine-tuning performance of some real-world datasets, pretraining the ICL-transformer on highly complex datasets is more important than pretraining on realistic datasets, although for zero-shot, we would always prefer the TabPFN dataset generator.
As we would like to have a single ICL-transformer that performs well across all real-world tabular datasets, we mix the TabPFN and forest dataset generators to pretrain TabForestPFN. This model has excellent fine-tuning performance on two benchmarks (Grinsztajn et al., 2022; McElfresh et al., 2023), matching the performance of either TabForest or TabPFN. At the same time, mixing in the forest dataset generator does not seem to harm the zero-shot performance at all. This makes TabForestPFN the preferred ICL-transformer over TabForest and TabPFN.
@@
Figure 1: Comparison of decision boundaries for the Electricity dataset (OpenML ID 44156). Axis represent features, colors are predicted class probabilities, and dots are test observations. Fine-tuned variants show a higher complexity score $V$ (see section 5.3) than zero-shot variants.
@@
In conclusion, fine-tuned ICL-transformers are highly effective tabular data classifiers, capable of creating complex decision boundaries. This new insight advances our understanding of tabular ICL-transformers and opens up new avenues for further research to enhance their performance. With further developments, we anticipate a significant shift in the field of tabular data, moving from tree-based methods towards ICL-transformers.
@@ 2 Related Works
There are three main branches of tools for tabular data classification: classical statistical methods like linear regression, K-nearest neighbors, Gaussian processes (Williams & Rasmussen, 1995), and support vector machines (Hearst et al., 1998); tree-based algorithms like XGBoost (Chen & Guestrin, 2016), CatBoost (Prokhorenkova et al., 2018), and LightGBM (Ke et al., 2017); and neural network-based methods such as the approach presented in this paper. There are several papers benchmarking the different methods (Gorishniy et al., 2021; Shwartz-Ziv & Armon, 2022; Grinsztajn et al., 2022; McElfresh et al., 2023; Zabërgja et al., 2024). Overall, tree-based methods stand at the top, with neural networks ranging from inferior to at best competitive.
Nonetheless, there have been numerous approaches that tackle tabular data classification with neural networks. First, we have the class of neural networks trained from scratch: training starts from random initialized weights and is only trained on the data at hand. Research has focused on architectures (Katzir et al., 2021; Somepalli et al., 2021; Arik & Pfister, 2021; Gorishniy et al., 2023; Huang et al., 2020; Chen et al., 2023a), embeddings (Ruiz et al., 2023; Gorishniy et al., 2022; Chen et al., 2023b), and regularization (Shavitt & Segal, 2018; Kadra et al., 2021).
In general, methods training from scratch can struggle because tabular datasets can be small. So, researchers have sought ways to use large volumes of tabular data or to change the training objective. Some employ self-supervised learning (Kossen et al., 2021; Yoon et al., 2020; Zhu et al., 2023; Bahri et al., 2022; Ucar et al., 2021; Sui et al., 2023), or closely related transfer learning techniques (Nam et al., 2023; Levin et al., 2023; Zhou et al., 2023). Others leverage pretrained LLMs or language data (Hegselmann et al., 2023; Zhang et al., 2023; Kim et al., 2024; Yan et al., 2024) to make predictions.
One of those related transfer learning methodologies is tabular in-context learning, a new field sparked by TabPFN (Hollmann et al., 2023). Currently, there is ongoing research on how to scale TabPFN to encompass more observations and features (Ma et al., 2023; Feuer et al., 2024; Thomas et al., 2024), as this architecture is limited by GPU memory. Our fine-tuning work can be seen as one approach to tackle this issue.
@@ 3 Preliminaries
In tabular classification, we are interested in predicting targets $y\in{\mathbb{N}}$ given features ${\bm{x}}\in{\mathbb{R}}^{d}$, where $d$ is the number of features. We predict $y$ using an in-context learning (ICL) transformer pretrained on a synthetic dataset. The in-context learning allows the transformer to predict targets based on other observations included in the forward pass. In our work, zero-shot refers to one forward pass through the ICL-transformer without any fine-tuning, while fine-tuning refers to one forward pass through the ICL-transformer after fine-tuning. Our work builds on TabPFN (Hollmann et al., 2023), so below we explain their dataset generator and their transformer architecture.
@@ 3.1 TabPFN Dataset Generator
The TabPFN authors create their own synthetic dataset using Bayesian Neural Networks (BNN) and Structural Causal Models (SCM). To construct a dataset, they first create a BNN or a SCM with random characteristics and with randomly initialized weights. Then they randomly draw an input ${\bm{X}}$ and pass it through the model to generate output ${\bm{y}}$. Their final dataset is given by $({\bm{X}},{\bm{y}})$. See their paper (Hollmann et al., 2023) for more details.
In their approach, they emphasize their ability to create realistic datasets, and even call their generator a “prior”. They chose SCMs specifically because it can capture real-world causal mechanisms. One other aspect they focus on is simplicity, biasing the generator towards less complex input-output relationships. Additionally, they ensure the inputs ${\bm{X}}$ have natural correlations by correlating the features blockwise, and they vary their feature importance by tuning the magnitude of weights belonging to different features. These methods suggest the authors believe creating realistic datasets is important for achieving good performance.
@@ 3.2 Architecture
In our work, we use the architecture from TabPFN, and make no changes to isolate the effect of the dataset generator. This ICL-transformer has as input the features ${\bm{X}}_{support}\in{\mathbb{R}}^{|S|\times d_{f}}$ and targets ${\bm{y}}_{support}\in{\mathbb{N}}^{|S|}$ from support set $S$ and features ${\bm{X}}_{query}\in{\mathbb{R}}^{|Q|\times d_{f}}$ from query set $Q$. The output is a prediction for ${\bm{y}}_{query}\in{\mathbb{R}}^{|Q|}$. The query set $Q$ represents the observations we want to predict, while the support set $S$ includes the observations we base our prediction on. This architecture accepts a fixed number of features $d_{f}$, see also the preprocessing discussed in Appendix A.2.
In this transformer, a token with dimension $d_{token}$ represents all features of a single observation. The creation of support tokens $H_{support}\in{\mathbb{R}}^{|S|\times d_{token}}$ and query tokens $H_{query}\in{\mathbb{R}}^{|Q|\times d_{token}}$ is given by equations 1 and 2.
@@ 4.1 Forest Dataset Generation
Table 1: Hyperparameters for the Forest Dataset Generator
Hyperparameter min max
base size 1024 1024
dataset size 128 1024
tree depth 1 25
number of features 3 100
number of classes 2 10
ratio of categorical features 0.0 1.0
Our goal is to create a simple dataset generator that produces datasets with complex patterns to train on, in contrast to the TabPFN (Hollmann et al., 2023) generator that aims to create realistic datasets. This forest dataset generator will better enable ICL-transformers to create complex decision boundaries. We base our dataset generator on decision trees, because of their ability to create highly complex decision boundaries with minimal computational cost. The idea is to overfit the decision tree to randomly generated features and targets. This fitted decision tree is then used as a data-generating process. See Algorithm 1 for the method and Figure 3 for examples of generated data.
@@
Our forest dataset generator allows datasets to vary in the number of classes, observations, numerical features, and categorical features. There are two parameters that contribute to the decision boundary complexity. The base size is the number of observations used to fit the decision tree; more observations means more places for the decision tree to split on. The tree depth determines how deep the decision tree will go before exiting the fitting algorithm, with higher depth leading to increased complexity.
For every new synthetically generated dataset, we uniformly draw the hyperparameters from the bounds shown in Table 1. Because both hyperparameters influence the complexity, we decided to keep the base size fixed. In the final step of the algorithm, the targets ${\bm{y}}_{2}$ are discretized by uniformly drawing bucket boundaries between 0.0 and 1.0, and assigning a class to each bucket, creating varying degrees of class imbalance.
Figure 3: Generated forest data. Every box is a generated dataset with its own classes (color) and features (axes). The data clouds look unrealistic: decision boundaries are always orthogonal, and there is no feature correlation. Generated with base size of 1024, dataset size of 1024, maximum tree depth between 1 and 25, two features, and between 2 and 10 number of classes.
@@ 4.2 Fine-tuning Procedure
In our work, we introduce fine-tuning to the tabular ICL-transformer. When fine-tuning, we like to draw support and query sets from our training data such that the performance generalizes to the test set. This requires careful consideration of dataset splitting. The benchmark datasets already provide us with a training-validation-test split. We use this validation dataset for early stopping.
Every gradient descent step, we randomly draw a 80/20 support and query split from the training dataset. For validation, we draw the support set from the training set and draw the query set from the validation set. Every validation sample is seen exactly once, while the support samples are randomly drawn with replacement. We use early stopping based on the validation loss, which is calculated after every fine-tuning step.
The early stopping technique can decide to stop fine-tuning immediately if the validation loss increases in the first step of training, which allows us to fall back on the zero-shot performance in case fine-tuning harms the performance. This is especially important when using very small datasets, as they are prone to overfitting. At the same time, fine-tuning can leverage all samples in training dataset, while zero-shot cannot use more samples than that fit on the GPU.
@@ 5 Experiments
In our experiments we consider five pre-trained models, each with a zero-shot and a fine-tuned version:
  • TabPFN (original) is the original implementation by the TabPFN (Hollmann et al., 2023) authors, fine-tuned by us. The weights are downloaded from their GitHub.
  • TabPFN (retrained) is our implementation of TabPFN, trained by us on the TabPFN-dataset.
@@
  • TabScratch is not pretrained. We include this as a baseline.
Comparing the behavior and performance allows us to understand the effect of the different synthetic datasets. Training and hyperparameter settings are given in appendix A.3, benchmarks used and results obtained are given below.